OpenStudy (anonymous):

Q tree casts a shadow of 28 feet. At the same time, a biologist who is 7 feet tall casts a shadow of 2 feet. How tall is the tree? A.8 feet B.14 feet C.78 feet D.98 feet

OpenStudy (anonymous):

Trick question :P

OpenStudy (anonymous):

How is it a trick question? The answer should be 98...

OpenStudy (anonymous):

Because, you didn't explain how you got 98 >:|

OpenStudy (anonymous):

Haha okay, 28/2 is 14, so 7*14 is 98! Voila!

OpenStudy (anonymous):

7/x=2/28 28*7=2x 156=2x 156/2=x 78=x

OpenStudy (anonymous):

I think its 78 because a 98 foot tall tree is insane. Im not sure though D:

OpenStudy (anonymous):

Ill try both xD

OpenStudy (anonymous):

Lol there are trees taller than that!

OpenStudy (anonymous):

your tree is 78 ft tall